painting if 50 feet by 30ft they want to put a border to be 25% of the painting what is the exact width of the border?
Answer by stanbon(75887)

painting of 50 feet by 30ft they want to put a border to be 25% of the painting what is the exact width of the border?
--------------------
If I read you correctly the area of the border is to be 25% of the area
of the painting.
--------------
The area of the painting is 50x30 = 1500 sq ft
25% of 1500 = 375 sq. ft.
===============================
Assume the width of the border is x ft.
At this point you should sketch an appropriate figure of the painting with
its border.
The width of the border is 50+2x ft.
The length of the border is 30+2x ft.
--------------------
So, the area of the picture INCLUDING its border is (50+2x)(30+2x) sq. ft.
------------------------------
EQUATION:
Entire area - picture area = 375 sq ft
(50+2x)(30+2x) - 1500 = 375
1500 + 160x + 4x^2 = 375
4x^2 + 160x = 1125
4x^2 + 160x - 1125 = 0
x=6.1007663... ft
=======================
Cheers,
Stan H.